OpenStudy (anonymous):

the line with a slope of 1/4 and y-intercept of -2

OpenStudy (skullpatrol):

What about it?

OpenStudy (jdoe0001):

hehe, I gather he means to find the equation of it =)

OpenStudy (jdoe0001):

y-intercept, means where the graph TOUCHES the y-axis, or INTERCEPTS it, like  notice that when that happens, what's the value of "x"? well, x =0, so if the y-intercept is at -2 and x =0, the point is ( 0, -2) to find the equation of a line with such point and slope you'd use the point-slope form and plug in the values, thus \(\bf \begin{array}{lllll} &x_1&y_1\\ &(0\quad ,&-2) \end{array} \\\quad \\ slope = m= \cfrac{1}{4} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (skullpatrol):

@jdoe0001 one of the many skills a student needs to learn is HOW to ask a question properly, right? We should not have to guess at his "intention," for productive learning to take place.

OpenStudy (anonymous):

thanks and sorry